OpenStudy (anonymous):

How many degrees are in each interior angle of a regular pentagon? 50 72 108 120

OpenStudy (lgbasallote):

use \[\large 180(n-2)\] a pentagon has 5 sides

OpenStudy (anonymous):

okay i thought so

OpenStudy (anonymous):

so then 180(5-2)?

OpenStudy (anonymous):

should be [180(5-2)]/5. 180(n-2) is total interior angle, so to get the value of each angle just divide it by n.

OpenStudy (lgbasallote):

yes @LovelyMegan

OpenStudy (anonymous):

@lgbasallote you have to divide it by n to get the value for each angle.

OpenStudy (anonymous):

ok im not getting any of the asnwer's

OpenStudy (anonymous):

answer's

OpenStudy (anonymous):

HEYYYYYYYY i got it its 108 right? plz tell me im right :)

OpenStudy (anonymous):

Formula to find the sum of all interior angles = 180(n-2) where n is the number of sides. First, use this formula to find the sum of all interior angles. In this case, it's a pentagon, so there will be 5 sides, whereby n=5. so sum of interior angles=180(5-2)=540. since it's a regular pentagon, all sides are of the same length, therefore all interior angles are of the same value. so to find the value for each angle, just take 540/5 which is 108 degrees.

OpenStudy (anonymous):

@LovelyMegan yes you're right. :)

OpenStudy (anonymous):

thanks